please answer areas with red x. i have tried these value positive and negative and they are not correct.GreenLawns provides a lawn fertilizing and weed control service. The company is adding a special aeration treatment as a low-
Answer #1

Note: All the parts have been done as shown in the attached images. In case of any doubt, please mention it in the comment. Thank you.

Page where Be Themen hemen advertising Sales -22_19M8-9RM +18R+33M dollars spent on radio adrest M- Thousands dollars pent diDate Page co We can Solve this using lagrange multipliers fCRM)= -28_12M298M +18R +33M R+M - 4 g (RM) Iccording to Jag rangeDate Page So -4R-9M + 18-1XX=0 and -24M - 9R +33 - 1x) - 0 -4R-9M +18 = -20-9R +33= from 7 & 1 -4R-9M + 18 = -20 9R + 33 5R +Date Put value of Mi Put value of M in TV D +R=4 of R=4 * Put value R & M in salu function is Sales = -2XC4) 12x 69_9x480 +18